# Better bounds on Grothendieck constants of finite orders

### v1

@@ IV.3 Exact upper bound on $K_{G}\left({3}\mapsto{2}\right)$
Let $P\in\mathrm{SDP}_{d}^{m_{1},m_{2}}$ with underlying vectors $a_{1},\ldots,a_{m_{1}},b_{1},\ldots,b_{m_{2}}\in S^{d-1}$ such that there exist $\eta_{A},\eta_{B}\in[0,1]$ for which $\eta_{A}S^{d-1}\subset\operatorname{conv}\{a_{x}\}_{x=1}^{m_{1}}$ and $\eta_{B}S^{d-1}\subset\operatorname{conv}\{b_{y}\}_{y=1}^{m_{2}}$. If $\alpha$ is such that $\alpha P\in\mathrm{SDP}_{n}^{m_{1},m_{2}}$, then
$K_{G}\left({d}\mapsto{n}\right)\leqslant\frac{1}{\alpha\,\eta_{A}\eta_{B}}.$ (14)
@@
Consider a matrix $N\in\mathrm{SDP}_{d}^{m_{1},m_{2}}$ and observe that the assumptions on $P$ imply that $\alpha\,\eta_{A}\eta_{B}\,N\in\mathrm{SDP}_{n}^{m_{1},m_{2}}$. Then, for any matrix $M$,
@@
$\alpha\,\eta_{A}\eta_{B}\leavevmode\nobreak\ \mathrm{SDP}_{d}\left({M}\right)=\max_{N\in\mathrm{SDP}_{d}^{m_{1},m_{2}}}\left\langle{M},{\alpha\,\eta_{A}\eta_{B}\,N}\right\rangle\leqslant\mathrm{SDP}_{n}\left({M}\right),$ (15)
so that Eq. 14 follows from the definition of $K_{G}\left({d}\mapsto{n}\right)$ in Eq. 5. ∎
@@
To obtain our bound on $K_{G}\left({3}\mapsto{2}\right)$, we start from the Gram matrix $P$ obtained from a centrosymmetric polyhedron with 912 vertices, that is, 406 lines. Note that this polyhedron is different from the one used in DIB+23 and features a slightly higher shrinking factor through the following construction: starting from an icosahedron, apply four times the transformation adding the normalised centers of each (triangular) facet. Taking $v_{0}=0.8962$ and running our Frank-Wolfe algorithm, we obtain a decomposition using 7886 extreme points of $\mathrm{SDP}_{2}^{406}$ and recovering $v_{0}P$ up to a Euclidean distance of $\varepsilon\approx 2.7\times 10^{-4}$. This algorithm is strictly identical to the one in DIB+23 except for the LMO, which corresponds to the case decribed in Section III.1 when $n=2$. Following (DIB+23, , Appendix D) we can make this decomposition purely analytical by noting that the unit ball for the Euclidean norm is contained in $\mathrm{SDP}_{1}$ and therefore in $\mathrm{SDP}_{2}$. Note that, contrary to DIB+23 where the extremal points of $\mathrm{SDP}_{1}$ were automatically analytical as they only have ±$\pm 1$ elements, an extra step is needed here to make numerical extremal points of $\mathrm{SDP}_{2}$ analytical. Here we follow a rationalisation procedure very similar to the one described in DIB+23 and we obtain a value $\alpha=v_{0}/(1+\varepsilon)$ that can be used in Eq. 14 to derive the bound
@@ VII Acknowledgements
The authors thank Péter Diviánszky for fruitful discussions and his help to run the branch-and-bound algorithm from DBV17 on asymmetric instances; Jim Reeds and Neil Sloane for sharing the details of the derivation of the value $45/31$ for the $E_{8}$ root system mentioned in FR94 ; Timotej Hrga, Nathan Krislock, Thi Thai Le, and Daniel Rehfeldt for trying to solve our symmetric instances with various solvers; Mathieu Besançon, Daniel Brosch, Wojciech Bruzda, Dardo Goyeneche, Volker Kaibel, David de Laat, Frank Valentin, and Karol Życzkowski for inspiring discussions. This research was partially supported by the DFG Cluster of Excellence MATH+ (EXC-2046/1, project id 390685689) funded by the Deutsche Forschungsgemeinschaft (DFG). T. V. acknowledges the support of the European Union (QuantERA eDICT) and the National Research, Development and Innovation Office NKFIH (Grants No. 2019-2.1.7-ERA-NET-2020-00003 and No. K145927).

### v2

@@ IV.3 Exact upper bound on $K_{G}\left({3}\mapsto{2}\right)$
Let $P\in\mathrm{SDP}_{d}^{p_{1},p_{2}}$ with underlying vectors $a_{1}^{P},\ldots,a_{p_{1}}^{P},b_{1}^{P},\ldots,b_{p_{2}}^{P}\in S^{d-1}$ such that
  • there exist $\eta_{1},\eta_{2}\in(0,1)$ for which $\eta_{1}S^{d-1}\subset\operatorname{conv}\{a_{i}^{P}\}_{i=1}^{p_{1}}$ and $\eta_{2}S^{d-1}\subset\operatorname{conv}\{b_{j}^{P}\}_{j=1}^{p_{2}}$,
  • there exists $\alpha\in(0,1)$ such that $\alpha P\in\mathrm{SDP}_{n}^{p_{1},p_{2}}$.
Then we have the following bound:
$K_{G}\left({d}\mapsto{n}\right)\leqslant\frac{1}{\alpha\,\eta_{1}\eta_{2}}.$ (14)
@@
Consider a matrix $M$ of size $m_{1}\times m_{2}$. Since
$\alpha\,\eta_{1}\eta_{2}\,\mathrm{SDP}_{d}\left({M}\right)=\alpha\,\eta_{1}\eta_{2}\max_{N\in\mathrm{SDP}_{d}^{m_{1},m_{2}}}\left\langle{M},{N}\right\rangle=\max_{N\in\mathrm{SDP}_{d}^{m_{1},m_{2}}}\left\langle{M},{\alpha\,\eta_{1}\eta_{2}\,N}\right\rangle,$ (15)
we focus on $\alpha\,\eta_{1}\eta_{2}\,N$ for $N\in\mathrm{SDP}_{d}^{m_{1},m_{2}}$ with underlying vectors $a_{1}^{N},\ldots,a_{m_{1}}^{N},b_{1}^{N},\ldots,b_{m_{2}}^{N}\in S^{d-1}$.
By the first assumption on $\{a_{i}^{P}\}_{i=1}^{p_{1}}$ and $\{b_{j}^{P}\}_{j=1}^{p_{2}}$, we know that we can write, for $x\in[m_{1}]$ and $y\in[m_{2}]$,
@@
$\eta_{1}a_{x}^{N}=\sum_{i=1}^{p_{1}}v_{i}^{(x)}a_{i}^{P}\quad\text{and}\quad\eta_{2}b_{y}^{N}=\sum_{j=1}^{p_{2}}w_{j}^{(y)}b_{j}^{P},\quad\text{so that}\quad\alpha\,\eta_{1}\eta_{2}\,N_{xy}=\alpha\sum_{i,j}v_{i}^{(x)}w_{j}^{(y)}\left\langle{a_{i}^{P}},{b_{j}^{P}}\right\rangle$ (16)
where $v_{i}^{(x)}\geqslant 0$, $w_{j}^{(y)}\geqslant 0$, $\sum_{i}v_{i}^{(x)}=1$, and $\sum_{j}w_{j}^{(y)}=1$.
Next we use the second assumption, which can be written as follows: there exist positive weights $\lambda_{k}$ summing up to one and vectors $\tilde{a}_{1}^{(k)},\ldots,\tilde{a}_{p_{1}}^{(k)},\tilde{b}_{1}^{(k)},\ldots,\tilde{b}_{p_{2}}^{(k)}\in S^{n-1}$ indexed by $k$ (in a finite set) such that $\alpha\langle{a_{i}^{P}},{b_{j}^{P}}\rangle=\sum_{k}\lambda_{k}\langle{\tilde{a}_{i}^{(k)}},{\tilde{b}_{j}^{(k)}}\rangle$ holds for all $i\in[p_{1}]$ and $j\in[p_{2}]$. This gives, for $x\in[m_{1}]$ and $y\in[m_{2}]$,
$\alpha\,\eta_{1}\eta_{2}\,N_{xy}=\sum_{i,j}v_{i}^{(x)}w_{j}^{(y)}\sum_{k}\lambda_{k}\left\langle{\tilde{a}_{i}^{(k)}},{\tilde{b}_{j}^{(k)}}\right\rangle=\sum_{k}\lambda_{k}\left\langle{\sum_{i}v_{i}^{(x)}\tilde{a}_{i}^{(k)}},{\sum_{j}w_{j}^{(y)}\tilde{a}_{j}^{(k)}}\right\rangle=\sum_{k}\lambda_{k}\left\langle{\tilde{a}_{x}^{(k)}},{\tilde{b}_{y}^{(k)}}\right\rangle,$ (17)
where we have defined the vectors $\tilde{a}_{x}^{(k)}=\sum_{i}v_{i}^{(x)}\tilde{a}_{i}^{(k)}$ and $\tilde{b}_{y}^{(k)}=\sum_{j}w_{j}^{(y)}\tilde{a}_{j}^{(k)}$ in the $n$-dimensional unit ball $B^{n}$.
Combining Eq. 15 with Eq. 17 and using the definition of $\mathrm{SDP}_{d}\left({M}\right)$ in Eq. 1, we obtain
$\alpha\,\eta_{1}\eta_{2}\,\mathrm{SDP}_{d}\left({M}\right)\leqslant\max\left\{\sum_{x=1}^{m_{1}}\sum_{y=1}^{m_{2}}M_{xy}\left\langle{\tilde{a}_{x}},{\tilde{b}_{y}}\right\rangle\leavevmode\nobreak\ \Big{|}\leavevmode\nobreak\ \forall x\in[m_{1}],\leavevmode\nobreak\ \tilde{a}_{x}\in B^{n},\leavevmode\nobreak\ \forall y\in[m_{2}],\leavevmode\nobreak\ \tilde{b}_{y}\in B^{n}\right\}\leqslant\mathrm{SDP}_{n}\left({M}\right),$ (18)
so that Eq. 14 follows from the definition of $K_{G}\left({d}\mapsto{n}\right)$. ∎
@@
To obtain our bound on $K_{G}\left({3}\mapsto{2}\right)$, we start from the Gram matrix $P$ obtained from a centrosymmetric polyhedron with 912 vertices, that is, 406 lines. Note that this polyhedron is different from the one used in DIB+23 and features a slightly higher shrinking factor through the following construction: starting from an icosahedron, apply four times the transformation adding the normalised centers of each (triangular) facet. Taking $v_{0}=0.8962$ and running our Frank-Wolfe algorithm, we obtain a decomposition using 7886 extreme points of $\mathrm{SDP}_{2}^{406}$ and recovering $v_{0}P$ up to a Euclidean distance of $\varepsilon\approx 2.7\times 10^{-4}$. This algorithm is strictly identical to the one in DIB+23 except for the LMO, which corresponds to the case decribed in Section III.1 when $n=2$. Following (DIB+23, , Appendix D) we can make this decomposition purely analytical by noting that the unit ball for the Euclidean norm is contained in $\mathrm{SDP}_{1}$ and therefore in $\mathrm{SDP}_{2}$. Note that, contrary to DIB+23 where the extremal points of $\mathrm{SDP}_{1}$ were automatically analytical as they only have ±$\pm 1$ elements, an extra step is needed here to make numerical extremal points of $\mathrm{SDP}_{2}$ analytical. Here we follow a rationalisation procedure very similar to the one described in (DIB+23, , Appendix A) and we obtain a value $\alpha=v_{0}/(1+\varepsilon)$ that can be used in Eq. 14 to derive the bound
@@ VII Acknowledgements
The authors thank Péter Diviánszky for fruitful discussions and his help to run the branch-and-bound algorithm from DBV17 on asymmetric instances; Jim Reeds and Neil Sloane for sharing the details of the derivation of the value $45/31$ for the $E_{8}$ root system mentioned in FR94 ; Timotej Hrga, Nathan Krislock, Thi Thai Le, and Daniel Rehfeldt for trying to solve our symmetric instances with various solvers; Fernando Mário de Oliveira Filho for asking details about the proof of Eq. 14, which helped us improve its presentation; Mathieu Besançon, Daniel Brosch, Wojciech Bruzda, Dardo Goyeneche, Volker Kaibel, David de Laat, Frank Valentin, and Karol Życzkowski for inspiring discussions. This research was partially supported by the DFG Cluster of Excellence MATH+ (EXC-2046/1, project id 390685689) funded by the Deutsche Forschungsgemeinschaft (DFG). T. V. acknowledges the support of the European Union (QuantERA eDICT) and the National Research, Development and Innovation Office NKFIH (Grants No. 2019-2.1.7-ERA-NET-2020-00003 and No. K145927).